### v1

# Autonomous Exploration and Semantic Updating of Large-Scale Indoor Environments with Mobile Robots

Abstract
We introduce a new robotic system that enables a mobile robot to autonomously explore an unknown environment, build a semantic map of the environment, and subsequently update the semantic map to reflect environment changes, such as location changes of objects. Our system leverages a LiDAR scanner for 2D occupancy grid mapping and an RGB-D camera for object perception. We introduce a semantic map representation that combines a 2D occupancy grid map for geometry, with a topological map for object semantics. This map representation enables us to effectively update the semantics by deleting or adding nodes to the topological map. Our system has been tested on a Fetch robot. The robot can semantically map a $93m\times 90m$ floor and update the semantic map once objects are moved in the environment.  [^1] Project page is available at: https://irvlutd.github.io/SemanticMapping
I Introduction
Autonomous exploration of unknown environments is an important problem in robotics. It has wide applications in robot search and rescue [1], environment surveillance [2] and service robots [3]. Early work on robot exploration focuses on building 2D occupancy grid maps [4] using LiDAR or sonar sensors [5, 6]. Later works consider using 3D octree maps [7] with RGB or RGB-D Simultaneous Localization and Mapping (SLAM) techniques such as ORB-SLAM [8] or KinectFusion [9]. Although these methods can build a geometric representation or map of an unknown environment, the map does not contain semantic information such as objects in the environment.
A common strategy to inject semantics into a geometric map is to leverage object detection or object segmentation using images, and then fuse the detected or segmented objects into the map. This process is commonly known as semantic mapping [10]. Objects can be represented using simple geometric primitives such as boxes or polygons [11, 12], or point clouds from depth cameras [13], or abstract nodes in a graph [14]. By combining robot exploration and semantic mapping, a semantic map of an environment can be built for downstream tasks such as object-goal navigation.
We observed that existing work on robot exploration and semantic mapping typically builds a map once and do not perform updating of the map. Consequently, these methods cannot handle environment changes such as objects being moved around in the environment. Environment changes are difficult for voxel-based or mesh-based mapping methods, since the voxel space or the 3D mesh needs to be updated.
Figure 1: Our system enables a mobile robot to explore a large-scale unknown environment, build a semantic map and update the map for environment changes.
In this work, we introduce a robotic system that is able to autonomously explore a large-scale indoor environment, build a semantic map of the environment online, and subsequently update the semantic map to account for environment changes. To the best of our knowledge, this is the first system that can perform all these tasks jointly on a physical robot. Specifically, our system first enables a mobile robot to explore an unknown environment and build a 2D occupancy grid map of the environment with a LiDAR scanner. Second, given the 2D map, the system plans a trajectory for the robot to traverse the whole environment. Third, the robot follows the planned trajectory and navigates in the environment. Meanwhile, during navigation, we apply open-vocabulary object detection and segmentation in real-time to images from the RGB-D camera on the robot. The detected objects are added to the 2D occupancy grid as nodes in a graph structure. Consequently, a semantic map of the environment can be constructed once the robot finishes the trajectory. Finally, in order to handle environment changes, the robot can visit the environment again and update the semantic map by deleting or adding object nodes online. Fig.  1 illustrates an overview of our system.
We have tested the system on the 4th floor of the Engineering and Computer Science South (ECSS) Building at the University of Texas at Dallas. Our Fetch robot was able to explore the whole floor and build a $93m\times 90m$ 2D occupancy grid map of the floor. In addition, we have shown that the robot can detect chairs, tables, and doors in the environment and build a semantic map for it. To test the effectiveness of the system on updating the semantic map, we intentionally moved some chairs and tables on the floor and let the robot traverse the environment again. We demonstrate that our system is able to update the semantic map to reflect changes in the environment.
II Related Work
II-A Robot Exploration without Semantics
Early works on robot exploration use LiDAR or sonar sensors to build 2D occupancy grid maps [4] of environments. A 2D occupancy grid contains cells that represent free space, obstacles, and unknown space. In 1997, Yamauchi introduced the concept of frontier in robot exploration [5], which built the foundation for frontier-based robot exploration algorithms [15, 6, 16, 17]. Generally, the frontiers on a map represent the boundaries between the unknown space and the free space. They are potential places for a robot to explore. Frontier-based exploration methods can differ in map representations, algorithms for detecting frontiers, and algorithms for selecting frontiers to visit. For example, [17] uses 3D octree maps [7], and the frontiers are 3D voxels in their approach. Most frontier-based exploration methods focus on building a metric map without semantics, such as a 2D occupancy grid map or a 3D octree map. In our method, we fuse semantic information into a 2D occupancy grid map during robot exploration.
II-B Robot Exploration with Semantics
Semantic information about objects in environments is important for robots to perform downstream tasks such as object-goal navigation [18] and off-road navigation [19]. Therefore, researchers incorporate semantics into robot exploration [20, 11, 21, 22]. Some methods simply utilize object segmentation or object detection to detect objects in the environment during exploration and then add semantic information to the map [20, 11]. Some methods consider semantics as another factor for the robot to explore [21, 22]. These methods jointly model geometry and semantics in robot exploration. In our method, we design a two-phase exploration strategy to deal with large-scale environments with large numbers of objects, where the first phase focuses on exploration of geometry to build a 2D occupancy grid map, and the second phase adds semantics into the 2D map.
II-C Semantic Mapping without Robot Exploration
In robot exploration, an exploration algorithm needs to be designed to control the movement of a robot, which also determines the movement of the sensors on the robot such as LiDAR scanners or RGB-D cameras. Therefore, the robot will have the ability to actively explore the environment. On the other hand, there are a large number of methods in the literature that focus on semantic mapping from offline data [23, 13, 24, 25, 26, 27]. These data can be collected by teleoperation of a mobile robot or using a hand-held camera, such as video sequences from a RGB-D camera. Generally speaking, these methods fuse semantic information about objects into a 3D map from an SLAM method. For example, SemanticFusion [13] combines ElasticFusion [28] for dense reconstruction with a neural network for semantic segmentation. DA-RNN [24] combines KinectFusion [9] and a recurrent neural network for video semantic segmentation. Recently, open-vocabulary object detectors such as GroundingDINO [29] have been utilized for semantic mapping. For example, ConceptGraphs [14] and its hierarchical version [30] build 3D scene graphs with open-vocabulary object detectors. Unlike these methods, the map in our method is built via robot exploration. Consequently, a robot can use the map to visit the environment again to refine and update the semantic information in the environment. Very few works have explored updating of the built semantic map except [12]. Our method improves over [12] in mapping larger environments autonomously with much more objects.
Figure 2: Illustration of our system for autonomous exploration, semantic mapping and map update.
III System
Our system is made up of two phases. In the initial phase, the mobile robot explores the environment and builds the occupancy map incrementally. In the subsequent phase, the robot leverages the built occupancy map to accurately localize itself within the environment and proceeds to construct the semantic map. An overview of the system architecture is presented in Fig. 2.
III-A Autonomous Exploration and Map Building
We have developed an environment exploration method based on the popular frontier exploration module [31], using a dynamic search window to efficiently navigate to frontiers in large, real-world environments.
Dynamic Window Frontier Exploration. We modified the this frontier exploration module [31] by dynamically adjusting the robot’s search area between local and global regions [^2] Code available at https://github.com/IRVLUTD/dynamic-window-frontier-exploration. This method identifies frontiers in the current search area and the robot navigates towards them to progressively map the environment. The search window is adjusted dynamically based on the density of the frontiers available around the robot. If the number of frontiers within the local search radius around the robot falls below a threshold, the search expands to a global radius. Setting this parameter to a large value expands the search region to the entire map. This allows the robot to move progressively ahead, reducing redundant coverage of the environment. The exploration continues until the number of frontiers falls below a global threshold or after a fixed time has elapsed, at which point the exploration terminates and saves the built map. Our improved exploration algorithm is integrated with the GMapping SLAM module [32, 33]. The approach is depicted in Fig. 2(a). During this process, the robot exploration trajectory is recorded which is later used to obtain a traversal plan. The saved occupancy map is then utilized to localize the robot during the semantic map construction.
III-B Autonomous Construction of Semantic Map
We utilize the saved occupancy map and the recorded robot exploration path to plan a trajectory to traverse the whole environment to construct the semantic map.
Figure 3: Environment traversal trajectory planning.
III-B1 Environment Traversing
To construct or update a semantic map, the robot needs to visit places in the environment and identify the objects. To optimize this travel, we first sample a subset of $n$ points $P=\{p_{1},p_{2},\dots,p_{n}\}$ along the robot exploration path (Fig. 3(a)), with a minimum of $2m$ separation between each pair of points (Fig. 3(b)). We then construct a graph $G(P,E)$ with the sampled way points as nodes and the Euclidean distance $d$ between them as corresponding edge weights $w_{ij}=d_{ij}=\|p_{i}-p_{j}\|$. This is depicted in Fig. 3(c). Finally, a traversal trajectory $T_{s}$ is obtained from $P$ using a greedy formulation of the Traveling Salesmen Problem (TSP) as explained in Algorithm 1. The generated trajectory is shown in Fig. 3(d).
Algorithm 1 Greedy Algorithm for TSP
1:Input: Sampled waypoints $P=\{p_{1},p_{2},\dots,p_{n}\}$, where $p_{1}$ is the starting waypoint
2:Output: A sequence of waypoints representing the TSP solution trajectory $T_{s}$
3:Initialize trajectory $T_{s}$ $\leftarrow[p_{1}]$
4:Initialize visited $\leftarrow\{p_{1}\}$
5:for $t\leftarrow 1$ to $n-1$ do
6:     Let $p_{\text{last}}\leftarrow$ last waypoint in the $T_{s}$
7:     $p_{\text{next}}\leftarrow\arg\min_{p_{j}\in P\setminus\text{visited}}w(p_{\text{last}},p_{j})$
8:     Append $p_{\text{next}}$ to $T_{s}$
9:     Add $p_{\text{next}}$ to visited
10:end for
11:Append $p_{1}$ to $T_{s}$$\triangleright$ Return to the starting waypoint
12:Return trajectory $T_{s}$
III-B2 Object Detection and Segmentation
During the traversal, object detection and segmentation is performed in real time to construct the semantic map. For real-time object detection, the system leverages GroundingDINO [29] to detect objects in the environment. It takes in the current RGB image from the robot’s camera, identifies objects, and generates labels and bounding boxes for each detection. These bounding boxes, along with the labels, are used as prompts for MobileSAM [34], a faster version implementation of Segment Anything [35], to generate the object segments.
After obtaining segments from the current frame and its corresponding depth image, we compute the 3D point cloud $\mathbf{P}_{\text{camera}}$ for each segment in camera frame by applying camera intrinsic parameters to the depth. These point clouds are then transformed into map frame of reference through the transformation matrices representing camera pose in robot base frame, $\mathbf{T}_{\text{camera}}\in\mathbb{SE}(3)$ and robot pose in map frame $\mathbf{T}_{\text{robot}}\in\mathbb{SE}(3)$:
$\mathbf{P}_{\text{map}}=\mathbf{T}_{\text{robot}}\cdot\mathbf{T}_{\text{camera}}\cdot\mathbf{P}_{\text{camera}}.$ (1)
Finally, the mean positions of the the point clouds are computed using $\mathbf{P}_{\text{mean}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{P}_{\text{map}}(i)$ for $m$ points in the point cloud as the 3D positions of the corresponding objects in the map frame .
III-B3 Semantic Map Representation and Construction
Representation. Our semantic map is a hybrid representation. At the lower level, we utilize a 2D occupancy grid map $M$ that captures the spatial information of the free space and obstacles in the environment. At the higher level, we construct a topological map $G_{T}(V)$, where $V=\{v_{1},v_{2},\dots,v_{n}\}$ are the nodes representing uniquely detected objects (Fig. 2(d)). Each node $v_{i}$ has the following attributes.
  • $v_{i}.category$: Object category label
  • $v_{i}.position$: Mean position of the object computed in the reference frame of the occupancy map. This attribute connects the topological map to the occupancy map in a hierarchical way.
  • $v_{i}.confidence$: confidence score of the detected object from object detection
This representation facilitates quicker object association and easy update of the nodes in the event of changes in environment.
Construction. To construct the semantic map, we first load the occupancy map $M$ built in Section III-A and localize the robot in it. Then an empty topological map $G_{T}(V)$, $V=\{\}$ is initialized. Additionally, for each object category (e.g., tables, chairs, doors), a category-specific object association distance threshold $\delta_{\{category\}}$ is defined. The robot then starts tracking the traversal trajectory generated in Section III-B1. Simultaneously, object detection and segmentation module is executed to detect objects in the field of view of the robot and compute their positions in the map $M$ frame of reference as illustrated in Fig. 4(a).
As depicted in Fig. 4(b), for every detected object in the robot field of view, its minimum distance to the existing nodes of the corresponding category is calculated. If this distance $<\delta_{\{category\}}$, then that particular object is considered to be already represented on the map and hence not added to $G_{T}$ again. If the minimum distance is greater than the threshold, then the object is added as a new node to $G_{T}$ as seen in Fig. 4(b). The result of this partial construction step is shown in Fig. 4(c). This process is continued until the robot completes the tracking of the traversal trajectory, with new nodes being added to $G_{T}$ as necessary. In this way, the semantic map is fully constructed, combining the geometric structure of the occupancy map $M$ with the layered semantic information of the topological map.
Figure 4: Adding objects to semantic map.
Updating. Once some changes are made in the environment, we proceed to update the semantic map to reflect these changes. This process is similar to the construction phase, but with additional validation steps to handle existing object removal, displacement and new object addition. First, the constructed semantic map is loaded, which acts as reference for the current state of the environment. As the robot tracks the traversal trajectory again, the detected objects in its current field of view are compared against those expected objects to be in the same field of view in the semantic map. If an expected object is no longer detected, it is assumed to have been removed or moved and the corresponding node is deleted from the map as shown in Fig. 5. Similarly, newly detected objects that do not match any existing nodes within distance threshold $\delta_{\{category\}}$ are added as new nodes, representing either new or relocated objects, updating the semantic map in real-time.
Figure 5: Removing objects from semantic map.
IV Experiments
The experiments are carried out in a large indoor environment measuring approximately $100m\times 100m$. It is made up of 18 corridors with a total traversal length of about $800m$. These corridors were mostly furnished with typical office furniture, including doors, tables, and chairs. Hence these 3 classes, i.e., table, door and chair are considered in the open vocabulary detection and segmentation module. We run the experiments on a Fetch mobile robot equipped with a LiDAR sensor and an RGB-D camera. Visual results can be found in Fig. 2.
Exploration and Mapping. To autonomously explore and map the environment, we integrated the mapping, exploration, and navigation modules.The Gmapping ROS package [33] was used to incrementally build a 0.1m/pixel occupancy grid map, using laser scan data from the Fetch robot. Simultaneously, we ran our custom-built Dynamic Window Frontier Exploration module to identify and select best valued frontier in the search window. The Move Base ROS package [36] was then used to navigate the robot to these frontiers. The robot’s maximum velocity was limited to 0.6 m/s with a 0.7m inflation radius to ensure safe navigation. After 150 minutes of exploration, the environment was fully mapped, and 2,250 trajectory points of exploration were recorded at 0.25Hz.
Environment Traversing. From the 2,250 robot poses recorded during the exploration phase, a subset of 130 poses are sampled with a minimum separation 2m distance between every pair of poses. A greedy travelling salesman problem was formulated with these poses and the sequence of sampled points for environment traversal is obtained as described in Sec. III-B1.
Real-time Semantic Map Construction. To enable real-time semantic mapping, the Fetch robot was equipped with a laptop featuring NVIDIA RTX 4090 GPU, running Ubuntu 20.04 and ROS Noetic. The laptop and Fetch robot were connected via Ethernet for subscribing and publishing the data through ROS topics.
Initially, the environment remained unchanged compared to the state during exploration. The system begins with built occupancy map. Localizing itself using AMCL ROS [36] it progressively builds the semantic map. Subsequently, semantic map updates were tested under the following distinct scenarios involving removal, addition, and moving of the objects in the environment.
Online Update I: 10 chairs were removed from a location while 5 chairs at another location have been moved significantly from their original positions.
Online Update II: 10 chairs were removed from a location. 5 chairs and a table have been added at another location and 5 more chairs at another different location. This update was conducted during the night.
IV-A Quantitative Analysis
Construction Update I Update II
Classes GT Map GT Map GT Map
Tables 45 45 45 39 45 37
Chairs 109 84 99 73 109 69
Doors 210 153 210 155 210 156
TABLE I: Semantic map object count across the construction and update stages. GT: Ground Truth, Map: Mapping result. The table compares the ground truth with the actual results across the construction, update I, and update II scenarios.
The robot traverses the environment, updating the semantic map based on the observed changes. The results of the semantic map construction and updates are presented in Table I. We can see that the detection of chairs and doors is relatively more challenging. During update I, the removal of 10 chairs after the construction phase is reflected in the decrease in the number of chairs from 84 to 73. However, in Update II, only 69 chairs were detected, differing from the expected result due to occlusions. Table II presents the breakdown of object detection performance across the three experimental phases. Doors have a relatively low mAP, mainly due to the it’s partial views while robot is traversing in narrow corridors. For chairs, the accuracy decreases during the update stages, indicating the challenges of occlusions.
Table Chair Door mAP
Scenario TP FP FN TP FP FN TP FP FN Table Chair Door
Construction 36 9 4 84 0 16 140 13 68 0.90 0.84 0.67
Update I 36 3 9 72 1 23 140 15 66 0.80 0.76 0.68
Update II 33 4 6 67 2 29 142 14 68 0.85 0.69 0.68
TABLE II: Object detection performance and mAP analysis. TP: True Positives, FP: False Positives, FN: False Negatives, mAP: Mean Average Precision for detected objects across the construction and update stages.
IV-B Qualitative Analysis
Figure 6: Examples of accurate detection
Figure 7: Representation of various challenging scenarios involved in object detection and semantic mapping.
Fig. 6 illustrates several examples of accurate object detection during our experiments for chairs, tables, and doors. In Fig. 7, we discuss representative challenging scenarios in object detection during robot navigation.
Occlusion Challenges. Objects like tables and chairs being in close proximity to each other, often results in occlusion. It leads to objects not being detected or segmented as a single object. This can be seen in Fig. 7(a) where table and chairs are segmented as table. In Fig. 7(c) table, chair, and a portion of the pillar are segmented together as table. In Fig. 7(b), a chair is partially occluded by the pillar. This can lead to incorrect object associations when the chair is detected from another view. In Fig. 7(e), the chairs in the background are completely occluded by the table and are not observed by the robot. Subsequently they are not associated with the semantic map.
Limitations in Constrained Spaces. The presence of constrained spaces like narrow corridors in the environment can hinder the robot’s ability to view the complete structure of objects such as doors as seen in Fig. 7(f). Many such doors are not detected, leading to reduction in the number of object nodes in the semantic map. This is evident in Table I.
Deviation from the Traversal Trajectory. In Fig. 7(d), large bin in the corridor occludes the objects behind it from the robot view, forces the robot to take a detour. This detour results in some objects being seen by the robot. Such scenarios greatly impact the quality of the semantic mapping, reducing the ability to accommodate environment changes.
Impact of Lighting. Varying lighting conditions can cause shadows and reflections. As seen in Fig. 7(g) and Fig. 7(h), these shadows on the floor are mistook for a door and a table, respectively, reducing the overall detection and semantic map accuracy. In Fig. 7(c), the overall environment brightness is reduced during the night, contributing to poor detection and segmentation.
Structural Misidentification. Walls resemble doors when viewed at certain angles, having similar rectangular structure. Some walls also have additional structures like support bars or cable cord covers, resembling large door handles. Hence, these walls are often misidentified as doors. One such case can be seen in Fig. 7(i).
V Conclusion and Future Work
We introduce a robotic system that enables a mobile robot to autonomously explore an unknown environment, recognize objects in the environment, and subsequently build a semantic map of the environment. The system utilizes a hybrid representation of the semantic map that consists of an occupancy grid map for geometry and a topological map for semantics. This representation enables us to update the semantic map efficiently to reflect the changes in the environment. Experiments conducted in a large indoor environment demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
Object perception during robot navigation still remains a challenge. Our future work will consider interactive object recognition, in which a robot can plan its actions to detect objects such as looking at some objects, moving towards an object to confirm its existence, and moving around to avoid occlusions.
Acknowledgement
This work was supported by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005 and the Sony Research Award Program. The work of T. Summers was supported by the United States Air Force Office of Scientific Research under Grant FA9550-23-1-0424 and the National Science Foundation under Grant ECCS-2047040. We would like to thank our colleague, Jishnu Jaykumar P, for his assistance during the experiments.

### v2

@@
 Our system leverages a LiDAR scanner for 2D occupancy grid mapping and an RGB-D camera for object perception.
-We introduce a semantic map representation that combines a 2D occupancy grid map for geometry, with a topological map for object semantics.
+We introduce a semantic map representation that combines a 2D occupancy grid map for geometry with a topological map for object semantics.
 This map representation enables us to effectively update the semantics by deleting or adding nodes to the topological map.
-Our system has been tested on a Fetch robot.
-The robot can semantically map a $93m\times 90m$ floor and update the semantic map once objects are moved in the environment.
-[^1] Project page is available at: https://irvlutd.github.io/SemanticMapping
+Our system has been tested on a Fetch robot, semantically mapping a $93m\times 90m$ and a $9m\times 13m$ indoor environment and updating their semantic maps once objects are moved in the environments.
+[^1] Project page with code is available at: https://irvlutd.github.io/SemanticMapping
 I Introduction
@@
 This process is commonly known as semantic mapping [10].
-Objects can be represented using simple geometric primitives such as boxes or polygons [11, 12], or point clouds from depth cameras [13], or abstract nodes in a graph [14].
-By combining robot exploration and semantic mapping, a semantic map of an environment can be built for downstream tasks such as object-goal navigation.
-We observed that existing work on robot exploration and semantic mapping typically builds a map once and do not perform updating of the map.
-Consequently, these methods cannot handle environment changes such as objects being moved around in the environment.
-Environment changes are difficult for voxel-based or mesh-based mapping methods, since the voxel space or the 3D mesh needs to be updated.
+Objects can be represented using simple geometric primitives such as boxes or polygons [11, 12], point clouds from depth cameras [13], or abstract nodes on a graph [14].
+By combining robot exploration and semantic mapping, a semantic map can be built for downstream tasks such as object-goal navigation [15, 16].
+We observed that existing work on robot exploration and semantic mapping typically builds a map once and does not perform updating of the map.
+Consequently, these methods cannot handle environment changes such as objects moving around in the environment.
+Environment changes are computationally intensive for voxel-based or mesh-based mapping methods, since the voxel space or the 3D mesh needs to be updated.
 Figure 1: Our system enables a mobile robot to explore a large-scale unknown environment, build a semantic map and update the map for environment changes.
 In this work, we introduce a robotic system that is able to autonomously explore a large-scale indoor environment, build a semantic map of the environment online, and subsequently update the semantic map to account for environment changes.
-To the best of our knowledge, this is the first system that can perform all these tasks jointly on a physical robot.
+To our knowledge, this is the first system that can perform all these tasks jointly and fully autonomously on a physical robot.
 Specifically, our system first enables a mobile robot to explore an unknown environment and build a 2D occupancy grid map of the environment with a LiDAR scanner.
-Second, given the 2D map, the system plans a trajectory for the robot to traverse the whole environment.
-Third, the robot follows the planned trajectory and navigates in the environment.
-Meanwhile, during navigation, we apply open-vocabulary object detection and segmentation in real-time to images from the RGB-D camera on the robot.
-The detected objects are added to the 2D occupancy grid as nodes in a graph structure.
+Second, given the 2D map, the system plans a trajectory for the robot to revisit the whole environment.
+Third, the robot follows the planned revisiting trajectory and navigates in the environment.
+During navigation, we apply an open-vocabulary object detection and an object segmentation model in real-time to images from the RGB-D camera on the robot.
+The detected objects are added to the topological map as nodes in a graph structure.
 Consequently, a semantic map of the environment can be constructed once the robot finishes the trajectory.
-Finally, in order to handle environment changes, the robot can visit the environment again and update the semantic map by deleting or adding object nodes online.
+Finally, to handle changes in the environment, the robot can visit the environment again and update the semantic map by removing or adding nodes to the map.
 Fig.
 1 illustrates an overview of our system.
-We have tested the system on the 4th floor of the Engineering and Computer Science South (ECSS) Building at the University of Texas at Dallas.
-Our Fetch robot was able to explore the whole floor and build a $93m\times 90m$ 2D occupancy grid map of the floor.
-In addition, we have shown that the robot can detect chairs, tables, and doors in the environment and build a semantic map for it.
-To test the effectiveness of the system on updating the semantic map, we intentionally moved some chairs and tables on the floor and let the robot traverse the environment again.
-We demonstrate that our system is able to update the semantic map to reflect changes in the environment.
+We validated our system in a large-scale $(\sim 8,500m^{2})$ and a medium-scale $(\sim 117m^{2})$ indoor environments, using a Fetch robot.
+The robot successfully explored, mapped, and updated its semantic map of the scene by detecting and adapting to changes such as moved furniture.
+Our main contributions in this work are as follows.
+  • End-to-End Autonomy: We integrate unseen environment exploration, mapping, revisiting, and semantic updating into one system that runs autonomously on a physical robot online.
+  • Practical Semantic Mapping: Using open-vocabulary object detection and segmentation, we build a hybrid representation of semantic map that efficiently scales to larger spaces.
+  • Continuous Environment Adaptation: Our system enables real-time semantic map updates to reflect dynamic changes in the environment.
+  • Physical System validation: We deploy and evaluate our system on a physical robot across large and medium-scale indoor environments, demonstrating the above contributions.
 II Related Work
@@
 These methods jointly model geometry and semantics in robot exploration.
-In our method, we design a two-phase exploration strategy to deal with large-scale environments with large numbers of objects, where the first phase focuses on exploration of geometry to build a 2D occupancy grid map, and the second phase adds semantics into the 2D map.
+In our method, we designed a two-phase exploration strategy to deal with large-scale environments with large numbers of objects, where the first phase focuses on exploration of geometry to build a 2D occupancy grid map, and the second phase adds semantics into the 2D map.
 II-C Semantic Mapping without Robot Exploration
 In robot exploration, an exploration algorithm needs to be designed to control the movement of a robot, which also determines the movement of the sensors on the robot such as LiDAR scanners or RGB-D cameras.
-Therefore, the robot will have the ability to actively explore the environment.
-On the other hand, there are a large number of methods in the literature that focus on semantic mapping from offline data [23, 13, 24, 25, 26, 27].
-These data can be collected by teleoperation of a mobile robot or using a hand-held camera, such as video sequences from a RGB-D camera.
+Thus, the robot will have the ability to actively explore the environment.
+There are a large number of methods in the literature that focus on semantic mapping from offline data [24, 13, 25, 26, 27, 28].
+This data can be collected by teleoperation of a mobile robot or using a hand-held camera, such as video sequences from a RGB-D camera.
+Therefore, these systems are not fully autonomous and depend on human control to explore and map the environments.
 Generally speaking, these methods fuse semantic information about objects into a 3D map from an SLAM method.
@@
 DA-RNN [25] combines KinectFusion [9] and a recurrent neural network for video semantic segmentation.
+[12] uses Mask R-CNN [30] for instance segmentation to construct and update a semantic map in real-time using a mobile manipulator system.
+However, it requires a fully explored environment map as a prerequisite to begin with.
+Hydra [31] builds and optimizes a 3D scene graph hierarchically from simulated and real-world datasets.
+Khronos [32] constructs dense spatio-temporal maps that reflect the 3D scene and also account for some dynamic environment changes while testing their method using a physical robot system.
+However, the robots are controlled manually during this process.
 Recently, open-vocabulary object detectors such as GroundingDINO [33] have been utilized for semantic mapping.
-For example, ConceptGraphs [14] and its hierarchical version [30] build 3D scene graphs with open-vocabulary object detectors.
-Unlike these methods, the map in our method is built via robot exploration.
-Consequently, a robot can use the map to visit the environment again to refine and update the semantic information in the environment.
-Very few works have explored updating of the built semantic map except [12].
-Our method improves over [12] in mapping larger environments autonomously with much more objects.
+For example, ConceptGraphs [14] and its hierarchical version [34] build 3D scene graphs with open-vocabulary object detectors based on a hand-held scanned dataset of the indoor environment.
+Although these works significantly advance in semantic mapping, they all share a common limitation, i.e., they are not fully autonomous.
+As shown in Table I, prior methods are based on manual control for environmental exploration.
+Some methods require pre-explored maps, and some do not perform semantic updating of environments.
+In contrast, our method initially builds the map via robot exploration.
+Consequently, the robot can use this map to revisit the environment in a planned and greedy manner, while continuously refining and updating the semantic information in the environment.
+This offers a fully autonomous approach to both mapping and semantic updating.
+Method Autonomous Exploration Mapping with Robot Semantic Mapping Semantic Updating
+SemanticFusion [13] × × ×
+DA-RNN [25] × × ×
+Hydra[31] × × ×
+ConceptGraphs[14] × × ×
+Dengler et al.[12] × ×
+Khronos[32] ×
+Ours
+TABLE I: Comparison of methods for autonomous exploration, semantic mapping and updating (reflecting environment changes).
 Figure 2: Illustration of our system for autonomous exploration, semantic mapping and map update.
 III System
-Our system is made up of two phases.
+Our system consists of two phases.
 In the initial phase, the mobile robot explores the environment and builds the occupancy map incrementally.
 In the subsequent phase, the robot leverages the built occupancy map to accurately localize itself within the environment and proceeds to construct the semantic map.
-An overview of the system architecture is presented in Fig.
+The motivation to separate the process into two stages, rather than performing simultaneously, is because of computational efficiency and robustness.
+Attempting to perform both tasks in a single stage could lead to resource constraints and reduced performance, especially for large-scale environments with a large number of objects.
+Additionally, in large environments, as the number of detected objects increases, the number of map corrections (like loop closures) applied to the detected objects increase, which slows down the mapping process.
+Evidently, this has also been mentioned in [32].
+Therefore, our two-phase approach is more efficient for semantic exploration and mapping of large-scale environments.
+An overview of the system is presented in Fig.
 2.
 III-A Autonomous Exploration and Map Building
-We have developed an environment exploration method based on the popular frontier exploration module [31], using a dynamic search window to efficiently navigate to frontiers in large, real-world environments.
+We have developed an environment exploration method based on the popular frontier exploration module [35], using a dynamic search window to efficiently navigate to frontiers in large real-world environments.
+When using [35] in large scale environments, as the map size increases, far-away frontiers become insignificant due to high cost, thus leaving a part of the map unexplored.
+Moreover, tuning the cost parameters in [35] is relatively hard and not intuitive.
+Figure 3: Environment traversal trajectory planning.
 Dynamic Window Frontier Exploration.
-We modified the this frontier exploration module [31] by dynamically adjusting the robot’s search area between local and global regions [^2] Code available at https://github.com/IRVLUTD/dynamic-window-frontier-exploration.
+We modified the frontier exploration module [35] by dynamically adjusting the robot’s search area between local and global regions [^2] Code available at https://github.com/IRVLUTD/dynamic-window-frontier-exploration.
 This method identifies frontiers in the current search area and the robot navigates towards them to progressively map the environment.
@@
 During this process, the robot exploration trajectory is recorded which is later used to obtain a traversal plan.
-The saved occupancy map is then utilized to localize the robot during the semantic map construction.
+The saved occupancy map is then utilized to localize the robot during the semantic map construction and update phases.
 III-B Autonomous Construction of Semantic Map
 We utilize the saved occupancy map and the recorded robot exploration path to plan a trajectory to traverse the whole environment to construct the semantic map.
-Figure 3: Environment traversal trajectory planning.
 III-B1 Environment Traversing
@@
 It takes in the current RGB image from the robot’s camera, identifies objects, and generates labels and bounding boxes for each detection.
-These bounding boxes, along with the labels, are used as prompts for MobileSAM [34], a faster version implementation of Segment Anything [35], to generate the object segments.
-After obtaining segments from the current frame and its corresponding depth image, we compute the 3D point cloud $\mathbf{P}_{\text{camera}}$ for each segment in camera frame by applying camera intrinsic parameters to the depth.
-These point clouds are then transformed into map frame of reference through the transformation matrices representing camera pose in robot base frame, $\mathbf{T}_{\text{camera}}\in\mathbb{SE}(3)$ and robot pose in map frame $\mathbf{T}_{\text{robot}}\in\mathbb{SE}(3)$:
+These bounding boxes are used as prompts for MobileSAM [38], a faster version of the Segment Anything Model (SAM) [39], to generate the object segmentation masks.
+After obtaining object segments of the current observation and its corresponding depth image, we compute the 3D point cloud $\mathbf{P}_{\text{camera}}$ for each segment in camera frame by applying camera intrinsic parameters to the depth.
+These points are then transformed to the map frame of reference through the transformation matrices representing camera pose in robot base frame, $\mathbf{T}_{\text{camera}}\in\mathbb{SE}(3)$ and robot pose in map frame $\mathbf{T}_{\text{robot}}\in\mathbb{SE}(3)$:
 $\mathbf{P}_{\text{map}}=\mathbf{T}_{\text{robot}}\cdot\mathbf{T}_{\text{camera}}\cdot\mathbf{P}_{\text{camera}}.$ (1)
-Finally, the mean positions of the the point clouds are computed using $\mathbf{P}_{\text{mean}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{P}_{\text{map}}(i)$ for $m$ points in the point cloud as the 3D positions of the corresponding objects in the map frame .
+Finally, the mean position of the point cloud is computed using $\mathbf{P}_{\text{mean}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{P}_{\text{map}}(i)$ for $m$ points in the point cloud as the 3D position of the corresponding object in the map frame.
 III-B3 Semantic Map Representation and Construction
-Representation.
 Our semantic map is a hybrid representation.
@@
 Each node $v_{i}$ has the following attributes.
-  • $v_{i}.category$: Object category label
-  • $v_{i}.position$: Mean position of the object computed in the reference frame of the occupancy map. This attribute connects the topological map to the occupancy map in a hierarchical way.
-  • $v_{i}.confidence$: confidence score of the detected object from object detection
-This representation facilitates quicker object association and easy update of the nodes in the event of changes in environment.
+  • $v_{i}.\text{id}$: A unique id of the object
+  • $v_{i}.\text{category}$: Object category label
+  • $v_{i}.\text{position}$: Mean position of the object computed in the reference frame of the occupancy map. This attribute connects the topological map to the occupancy map in a hierarchical way.
+  • $v_{i}.\text{confidence}$: the confidence score of the detected object from object detection
+This representation facilitates fast object association and easy update of the nodes in the event of changes in the environment.
 Construction.
-To construct the semantic map, we first load the occupancy map $M$ built in Section III-A and localize the robot in it.
+To construct the semantic map, we first load the occupancy map $M$ built in Section III-A and localize the robot using it.
 Then an empty topological map $G_{T}(V)$, $V=\{\}$ is initialized.
-Additionally, for each object category (e.g., tables, chairs, doors), a category-specific object association distance threshold $\delta_{\{category\}}$ is defined.
+Additionally, for each object category (e.g., tables, chairs, doors), a category-specific object association distance threshold $\delta_{\{\text{category}\}}$ is defined.
 The robot then starts tracking the traversal trajectory generated in Section III-B1.
@@
 4(b), for every detected object in the robot field of view, its minimum distance to the existing nodes of the corresponding category is calculated.
-If this distance $<\delta_{\{category\}}$, then that particular object is considered to be already represented on the map and hence not added to $G_{T}$ again.
-If the minimum distance is greater than the threshold, then the object is added as a new node to $G_{T}$ as seen in Fig.
+If this distance is small than $\delta_{\{\text{category}\}}$, then that particular object is considered to be already represented on the map and hence not added to the topological map $G_{T}$ again.
+If the minimum distance is greater than the threshold, then the object is added as a new node to $G_{T}$ as illustrated in Fig.
 4(b).
@@
 4(c).
-This process is continued until the robot completes the tracking of the traversal trajectory, with new nodes being added to $G_{T}$ as necessary.
-In this way, the semantic map is fully constructed, combining the geometric structure of the occupancy map $M$ with the layered semantic information of the topological map.
+This process is continued until the robot completes the tracking of the traversal trajectory, with new nodes being added to $G_{T}$ after object association.
+In this way, the semantic map is fully constructed, combining the geometric structure of the occupancy map with the layered semantics of the topological map.
 Figure 4: Adding objects to semantic map.
 Updating.
-Once some changes are made in the environment, we proceed to update the semantic map to reflect these changes.
-This process is similar to the construction phase, but with additional validation steps to handle existing object removal, displacement and new object addition.
-First, the constructed semantic map is loaded, which acts as reference for the current state of the environment.
-As the robot tracks the traversal trajectory again, the detected objects in its current field of view are compared against those expected objects to be in the same field of view in the semantic map.
+Once some changes are made to the environment, we update the semantic map to reflect these changes.
+This process is similar to the construction phase, but with additional validation steps to handle existing object removal, displacement, and new object addition.
+First, the constructed semantic map is loaded, which acts as a reference for the current state of the environment.
+As the robot tracks the traversal trajectory again, the detected objects in its current field of view are compared against those objects that are expected to be in the same field of view in the semantic map.
 If an expected object is no longer detected, it is assumed to have been removed or moved and the corresponding node is deleted from the map as shown in Fig.
 5.
-Similarly, newly detected objects that do not match any existing nodes within distance threshold $\delta_{\{category\}}$ are added as new nodes, representing either new or relocated objects, updating the semantic map in real-time.
+Similarly, newly detected objects that do not match any existing nodes within the distance threshold $\delta_{\{\text{category}\}}$ are added as new nodes, representing new or relocated objects.
+This process updates the semantic map in real-time.
 Figure 5: Removing objects from semantic map.
 IV Experiments
-The experiments are carried out in a large indoor environment measuring approximately $100m\times 100m$.
-It is made up of 18 corridors with a total traversal length of about $800m$.
+We have conducted experiments in two different environments with different object categories and evaluate the performance of our system in those environments.
+Environment A.
+The fourth floor of the ECSS building at UT Dallas is a large indoor environment measuring approximately $93m\times 90m$.
+It consists of 18 corridors with a total traversal length of approximately $800m$.
 These corridors were mostly furnished with typical office furniture, including doors, tables, and chairs.
-Hence these 3 classes, i.e., table, door and chair are considered in the open vocabulary detection and segmentation module.
-We run the experiments on a Fetch mobile robot equipped with a LiDAR sensor and an RGB-D camera.
-Visual results can be found in Fig.
-2.
+Therefore, we only consider these 3 classes, i.e., table, door and chair for semantic mapping of this environment.
+Environment B.
+This environment is a laboratory environment with a reduced size measuring $9m\times 13m$.
+This space included a more diverse set of objects such as televisions, trash bins, umbrellas, chairs, cabinets and persons.
+Robot. We run the experiments on a Fetch mobile robot equipped with a LiDAR sensor and an RGB-D camera as depicted in Fig. 6.
+Figure 6: The Fetch robot maps the environment using a $220^{o}$ laser scanner of $25m$ range, while leveraging depth observations for obstacle avoidance.
+Its RGB-D camera operates at 30fps, enabling real-time object detection and segmentation for semantic mapping running on the connected GPU laptop.
 Exploration and Mapping.
-To autonomously explore and map the environment, we integrated the mapping, exploration, and navigation modules.The Gmapping ROS package [33] was used to incrementally build a 0.1m/pixel occupancy grid map, using laser scan data from the Fetch robot.
+To autonomously explore and map the environments, we integrate the mapping, exploration, and navigation modules.
+The Gmapping ROS package [37] was used to incrementally build a 0.1m/pixel occupancy grid map, using laser scan data from the Fetch robot.
 Simultaneously, we ran our custom-built Dynamic Window Frontier Exploration module to identify and select best valued frontier in the search window.
-The Move Base ROS package [36] was then used to navigate the robot to these frontiers.
-The robot’s maximum velocity was limited to 0.6 m/s with a 0.7m inflation radius to ensure safe navigation.
-After 150 minutes of exploration, the environment was fully mapped, and 2,250 trajectory points of exploration were recorded at 0.25Hz.
+The move_base ROS package [40] guided the robot to these frontiers by combining a global planner for long term path planning and a local planner that dynamically avoids obstacles not seen during global planning.
+This ensures safe navigation when any new objects are added or moved around in the environment.
+The robot’s maximum velocity was limited to 0.6m/s, with a 0.7m inflation radius to prevent it from traveling too close to the obstacles.
+In Environment A, after 150 minutes of exploration, which also included several stops due to close proximity of persons in narrow corridors, the environment was fully mapped, and 2,250 trajectory points were recorded at 0.25Hz.
+In Environment B, exploration was completed in approximately 4 minutes, with 212 points recorded at 1Hz.
 Environment Traversing.
-From the 2,250 robot poses recorded during the exploration phase, a subset of 130 poses are sampled with a minimum separation 2m distance between every pair of poses.
-A greedy travelling salesman problem was formulated with these poses and the sequence of sampled points for environment traversal is obtained as described in Sec.
-III-B1.
-Real-time Semantic Map Construction.
+The greedy TSP described in Sect.
+III-B1 is formulated using the poses recorded during the exploration, resulting in traversal trajectories of 130 poses for Environment A and 15 poses for Environment B.
+In particular, the time to revisit is significantly reduced compared to the exploration phase.
+Revisiting took approximately 35 minutes for Environment A and 2 minutes for Environment B.
+This highlights the efficiency gained through this traversal optimization.
+Real-time Construction and Updating of Semantic Maps.
 To enable real-time semantic mapping, the Fetch robot was equipped with a laptop featuring NVIDIA RTX 4090 GPU, running Ubuntu 20.04 and ROS Noetic.
 The laptop and Fetch robot were connected via Ethernet for subscribing and publishing the data through ROS topics.
-Initially, the environment remained unchanged compared to the state during exploration.
-The system begins with built occupancy map.
-Localizing itself using AMCL ROS [36] it progressively builds the semantic map.
-Subsequently, semantic map updates were tested under the following distinct scenarios involving removal, addition, and moving of the objects in the environment.
-Online Update I: 10 chairs were removed from a location while 5 chairs at another location have been moved significantly from their original positions.
-Online Update II: 10 chairs were removed from a location.
-5 chairs and a table have been added at another location and 5 more chairs at another different location.
-This update was conducted during the night.
+Initially, the environment remained unchanged compared to its state during exploration.
+The system then begins with the built occupancy map.
+After localizing itself using AMCL ROS [40], the robot progressively builds the semantic map.
+Subsequently, semantic map updates were tested by removal, addition, and/or moving the objects in the environments.
+Visual results of environment A can be found in Fig.
+2.
+More details of it can be found in the supplementary video.
+Fig.
+7 depicts some of the changes in Environment B, reflected in the semantic updating stages.
+Figure 7: Visualization of some selected environmental changes in Environment B across two update stages.
+In Update stage I, 2 trash bins are removed.
+After that in update stage II, 2 chairs were removed.
+When robot revisited, these changes are reflected in the updated semantic map.
+Env Class Construction Update I Update II
+GT Map GT Map GT Map
+A Table 45 45 45 39 45* 37
+Chair 109 84 99 73 109* 69
+Door 170 153 170 155 170 156
+B Person 2 2 2* 2 2 2
+Trash Bin 2 2 0 0 0 1
+Umbrella 1 1 0 0 0 0
+Chair 3 3 3 3 1* 1
+Television 1 2 1 1 1 1
+Cabinet 3 2 3 1 3 2
+TABLE II: Semantic map object count across the construction and update stages for A & B Environments.
+GT: Ground Truth, Map: no.
+of objects mapped.
+The table compares the ground truth with the actual results across the construction, update I, and update II scenarios.
+* indicates that some objects have been relocated from their earlier position.
 IV-A Quantitative Analysis
-Construction Update I Update II
-Classes GT Map GT Map GT Map
-Tables 45 45 45 39 45 37
-Chairs 109 84 99 73 109 69
-Doors 210 153 210 155 210 156
-TABLE I: Semantic map object count across the construction and update stages.
-GT: Ground Truth, Map: Mapping result.
-The table compares the ground truth with the actual results across the construction, update I, and update II scenarios.
-The robot traverses the environment, updating the semantic map based on the observed changes.
-The results of the semantic map construction and updates are presented in Table I.
-We can see that the detection of chairs and doors is relatively more challenging.
-During update I, the removal of 10 chairs after the construction phase is reflected in the decrease in the number of chairs from 84 to 73.
-However, in Update II, only 69 chairs were detected, differing from the expected result due to occlusions.
-Table II presents the breakdown of object detection performance across the three experimental phases.
-Doors have a relatively low mAP, mainly due to the it’s partial views while robot is traversing in narrow corridors.
-For chairs, the accuracy decreases during the update stages, indicating the challenges of occlusions.
-Table Chair Door mAP
-Scenario TP FP FN TP FP FN TP FP FN Table Chair Door
-Construction 36 9 4 84 0 16 140 13 68 0.90 0.84 0.67
-Update I 36 3 9 72 1 23 140 15 66 0.80 0.76 0.68
-Update II 33 4 6 67 2 29 142 14 68 0.85 0.69 0.68
-TABLE II: Object detection performance and mAP analysis.
-TP: True Positives, FP: False Positives, FN: False Negatives, mAP: Mean Average Precision for detected objects across the construction and update stages.
+Table II summarizes the number of objects detected in both Environment A and Environment B during the Construction, Update I and Update II phases.
+In Environment A, we can see that while doors and tables were stably detected, the detection of chairs was relatively more challenging.
+During update I, the removal of 10 chairs after the construction phase is reflected in the decrease in the number of chairs detected from 84 to 73.
+However, in Update II, only 69 chairs were detected, with false negatives increasing to 29 as shown in Table III.
+This was mainly because some relocated chairs were placed in the path of the robot traversal trajectory.
+As a result, robot took a detour to avoid hitting them, causing it to miss some chairs that were not relocated and previously detected.
+Moreover, people occupying chairs occluded them, increasing false negatives.
+In contrast, Environment B has very stable detections, with many classes having 100% precision.
+However, some categories like cabinets that are dark colored, are challenging to detect in nominal lightning, resulting in them not being detected.
 IV-B Qualitative Analysis
-Figure 6: Examples of accurate detection
+Figure 8: Examples of accurate detection and segmentation of chairs, tables and doors in Environment A.
+Class Construction Update I Update II
+TP FP FN Precision TP FP FN Precision TP FP FN Precision
+Environment A
+Table 36 9 4 0.9 36 3 9 0.8 33 4 6 0.85
+Chair 84 0 16 0.84 72 1 23 0.76 67 2 29 0.69
+Door 140 13 28 0.83 140 15 26 0.84 142 14 28 0.84
+Environment B
+Person 2 0 0 1 2 0 0 1 2 0 0 1
+Trash Bin 2 0 0 1 0 0 0 - 0 1 0 0
+Umbrella 1 0 0 1 0 0 0 - 0 0 0 -
+Chair 3 0 0 1 3 0 0 1 1 0 0 1
+Television 1 1 0 0.5 1 0 0 1 1 0 0 1
+Cabinet 2 0 1 0.67 1 0 2 0.33 2 0 1 0.67
+TABLE III: Semantic map object count across the construction and update stages for A & B Environments.
+TP: True Positives, FP: False Positives, FN: False Negatives, Precision: Precision for detected objects across the construction and update stages.
+In Fig. 9, we show several challenging scenarios in object detection and segmentation, while robot is navigating.
+Occlusion Challenges.
+Close proximity of objects such as tables and chairs causes occlusion and leads to missed detections or incorrect segmentation, as seen in Fig.
+9(a).
+In Fig.
+9(c) table, chair, and a portion of the pillar are segmented together as table.
+The scenario seen in Fig.
+9(b), where a chair is partially occluded by the pillar, leads to incorrect object associations when the chair is detected again at another view.
+Detection Threshold.
+Lowering the detection threshold increases false positives, while a higher threshold may miss objects.
+In Fig.
+9(e), chairs in the foreground are undetected at a 0.8 threshold.
+careful choice of threshold for object class is important for accurate semantic mapping.
+Limitations in Constrained Spaces.
+Narrow corridors limit the robot’s view of objects like doors, as seen in Fig.
+9(f).
+Several of such doors are not detected, leading to a reduction in the number of object nodes in the semantic map.
+This is evident in Table II.
 Figure 9: Representation of various challenging scenarios involved in object detection and semantic mapping.
-Fig.
-6 illustrates several examples of accurate object detection during our experiments for chairs, tables, and doors.
-In Fig.
-7, we discuss representative challenging scenarios in object detection during robot navigation.
-Occlusion Challenges.
-Objects like tables and chairs being in close proximity to each other, often results in occlusion.
-It leads to objects not being detected or segmented as a single object.
-This can be seen in Fig.
-7(a) where table and chairs are segmented as table.
-In Fig.
-7(c) table, chair, and a portion of the pillar are segmented together as table.
-In Fig.
-7(b), a chair is partially occluded by the pillar.
-This can lead to incorrect object associations when the chair is detected from another view.
-In Fig.
-7(e), the chairs in the background are completely occluded by the table and are not observed by the robot.
-Subsequently they are not associated with the semantic map.
-Limitations in Constrained Spaces.
-The presence of constrained spaces like narrow corridors in the environment can hinder the robot’s ability to view the complete structure of objects such as doors as seen in Fig.
-7(f).
-Many such doors are not detected, leading to reduction in the number of object nodes in the semantic map.
-This is evident in Table I.
 Deviation from the Traversal Trajectory.
 In Fig.
-7(d), large bin in the corridor occludes the objects behind it from the robot view, forces the robot to take a detour.
-This detour results in some objects being seen by the robot.
-Such scenarios greatly impact the quality of the semantic mapping, reducing the ability to accommodate environment changes.
+9(d), a large bin in the corridor occluded objects, forcing the robot to take a detour.
+This detour resulted in the robot not seeing some objects.
+Such scenarios impact the quality of the semantic mapping, reducing the ability to accommodate environmental changes.
 Impact of Lighting.
@@
 As seen in Fig.
-7(g) and Fig.
-7(h), these shadows on the floor are mistook for a door and a table, respectively, reducing the overall detection and semantic map accuracy.
+9(g) and Fig.
+9(h), these shadows on the floor were mistook for a door and a table, respectively, reducing the overall detection and semantic map accuracy.
 In Fig.
-7(c), the overall environment brightness is reduced during the night, contributing to poor detection and segmentation.
+9(c), reduced brightness at night degraded detection and segmentation.
 Structural Misidentification.
-Walls resemble doors when viewed at certain angles, having similar rectangular structure.
-Some walls also have additional structures like support bars or cable cord covers, resembling large door handles.
-Hence, these walls are often misidentified as doors.
-One such case can be seen in Fig.
-7(i).
+Walls resembling doors due to shape or added features (e.g., support bars) often cause misidentification, as seen in Fig.
+9(i).
+This scenario also occurred in Environment B, update II phase, where a cardboard box was identified as a trash bin, leading to a false positive.
 V Conclusion and Future Work
@@
 This representation enables us to update the semantic map efficiently to reflect the changes in the environment.
-Experiments conducted in a large indoor environment demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
+Experiments conducted on a mobile manipulatory system in a large- and medium-scale indoor environments demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
 Object perception during robot navigation still remains a challenge.
-Our future work will consider interactive object recognition, in which a robot can plan its actions to detect objects such as looking at some objects, moving towards an object to confirm its existence, and moving around to avoid occlusions.
-Acknowledgement
+Our future work will consider interactive object recognition, in which a robot can plan its actions to detect objects such as looking at some objects, moving towards an object to confirm its existence for a robust object association, and dynamic planning of traversal to avoid occlusions.
+Acknowledgment
 This work was supported by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005 and the Sony Research Award Program.

### v3

@@
-Autonomous Exploration and Semantic Updating of Large-Scale Indoor Environments with Mobile Robots
+A Modular Robotic System for Autonomous Exploration and Semantic Updating in Large-Scale Indoor Environments
 Abstract
-We introduce a new robotic system that enables a mobile robot to autonomously explore an unknown environment, build a semantic map of the environment, and subsequently update the semantic map to reflect environment changes, such as location changes of objects.
-Our system leverages a LiDAR scanner for 2D occupancy grid mapping and an RGB-D camera for object perception.
-We introduce a semantic map representation that combines a 2D occupancy grid map for geometry with a topological map for object semantics.
-This map representation enables us to effectively update the semantics by deleting or adding nodes to the topological map.
-Our system has been tested on a Fetch robot, semantically mapping a $93m\times 90m$ and a $9m\times 13m$ indoor environment and updating their semantic maps once objects are moved in the environments.
+We present a modular robotic system for autonomous exploration and semantic updating of large-scale unknown environments.
+Our approach enables a mobile robot to build, revisit, and update a hybrid semantic map that integrates a 2D occupancy grid for geometry with a topological graph for object semantics.
+Unlike prior methods that rely on manual teleoperation or precollected datasets, our two-phase approach achieves end-to-end autonomy: first, a modified frontier-based exploration algorithm with dynamic search windows constructs a geometric map; second, using a greedy trajectory planner, environments are revisited, and object semantics are updated using open-vocabulary object detection and segmentation.
+This modular system, compatible with any metric SLAM framework, supports continuous operation by efficiently updating the semantic graph to reflect short-term and long-term changes such as object relocation, removal, or addition.
+We validate the approach on a Fetch robot in real-world indoor environments of approximately 8,5008,500m2 and 117117m2, demonstrating robust and scalable semantic mapping and continuous adaptation, marking a fully autonomous integration of exploration, mapping, and semantic updating on a physical robot.
 [^1] Project page with code is available at: https://irvlutd.github.io/SemanticMapping
@@
 In this work, we introduce a robotic system that is able to autonomously explore a large-scale indoor environment, build a semantic map of the environment online, and subsequently update the semantic map to account for environment changes.
-To our knowledge, this is the first system that can perform all these tasks jointly and fully autonomously on a physical robot.
+Our system represents a significant step forward in integrating autonomous exploration, long-term semantic mapping, and updating on a physical robot, addressing limitations of prior methods that rely on manual control or posed datasets.
 Specifically, our system first enables a mobile robot to explore an unknown environment and build a 2D occupancy grid map of the environment with a LiDAR scanner.
@@
 1 illustrates an overview of our system.
-We validated our system in a large-scale $(\sim 8,500m^{2})$ and a medium-scale $(\sim 117m^{2})$ indoor environments, using a Fetch robot.
-The robot successfully explored, mapped, and updated its semantic map of the scene by detecting and adapting to changes such as moved furniture.
+We validated our system in a large-scale 8,5008,500m2 and a medium-scale 117117m2 indoor environments, using a Fetch robot.
+The robot successfully explored, and updated the semantic map by detecting changes such as moved furniture.
 Our main contributions in this work are as follows.
-  • End-to-End Autonomy: We integrate unseen environment exploration, mapping, revisiting, and semantic updating into one system that runs autonomously on a physical robot online.
-  • Practical Semantic Mapping: Using open-vocabulary object detection and segmentation, we build a hybrid representation of semantic map that efficiently scales to larger spaces.
-  • Continuous Environment Adaptation: Our system enables real-time semantic map updates to reflect dynamic changes in the environment.
-  • Physical System validation: We deploy and evaluate our system on a physical robot across large and medium-scale indoor environments, demonstrating the above contributions.
+  • End-to-End Autonomy: We integrate unseen environment exploration, mapping, revisiting, and semantic updating into a modular system that operates fully autonomously on a physical robot in real-time.
+  • Lightweight Object Association: An efficient association strategy that enables fast and scalable semantic map updates.
+  • Short and Long-Term Semantic Updating: Our semantic mapping approach captures changes during the current visit as well as changes observed across multiple revisits, enabling the map to reflect both immediate and accumulated environmental updates.
+  • Real-World Validation at Scale: We deploy and evaluate our system on a physical robot in large- and medium-scale indoor environments, demonstrating the above contributions.
 II Related Work
@@
 In our method, we designed a two-phase exploration strategy to deal with large-scale environments with large numbers of objects, where the first phase focuses on exploration of geometry to build a 2D occupancy grid map, and the second phase adds semantics into the 2D map.
+Method Autonomous Exploration Metric Mapping with Robot Semantic Mapping Semantic Updating
+SemanticFusion[13] × × ×
+DA-RNN [24] × × ×
+Hydra[25] × × ×
+ConceptGraphs[14] × × ×
+Dengler et al.[12] × ×
+Khronos[26] ×
+Ours
+TABLE I: Comparison of methods for autonomous exploration, metric mapping with a physical robot, semantic mapping and updating (reflecting environment changes).
 II-C Semantic Mapping without Robot Exploration
@@
 Thus, the robot will have the ability to actively explore the environment.
-There are a large number of methods in the literature that focus on semantic mapping from offline data [24, 13, 25, 26, 27, 28].
+There are a large number of methods in the literature that focus on semantic mapping of offline data [27, 13, 24, 28, 29, 30].
 This data can be collected by teleoperation of a mobile robot or using a hand-held camera, such as video sequences from a RGB-D camera.
-Therefore, these systems are not fully autonomous and depend on human control to explore and map the environments.
+Therefore, these systems are not fully autonomous and depend on human control to explore and map environments.
 Generally speaking, these methods fuse semantic information about objects into a 3D map from an SLAM method.
@@
 Khronos [26] constructs dense spatio-temporal maps that reflect the 3D scene and also account for some dynamic environment changes while testing their method using a physical robot system.
-However, the robots are controlled manually during this process.
+However, robots are manually controlled during this process.
 Recently, open-vocabulary object detectors such as GroundingDINO [33] have been utilized for semantic mapping.
-For example, ConceptGraphs [14] and its hierarchical version [34] build 3D scene graphs with open-vocabulary object detectors based on a hand-held scanned dataset of the indoor environment.
+For example, ConceptGraphs [14] and its hierarchical version [34] build 3D scene graphs with open-vocabulary object detectors based on a hand-h eld scanned dataset of the indoor environment, having ground truth camera poses.
 Although these works significantly advance in semantic mapping, they all share a common limitation, i.e., they are not fully autonomous.
@@
 Some methods require pre-explored maps, and some do not perform semantic updating of environments.
-In contrast, our method initially builds the map via robot exploration.
+In contrast, our method initially builds the map through robot exploration.
+Figure 2: Illustration of our system for autonomous exploration, semantic mapping and map update.
 Consequently, the robot can use this map to revisit the environment in a planned and greedy manner, while continuously refining and updating the semantic information in the environment.
 This offers a fully autonomous approach to both mapping and semantic updating.
-Method Autonomous Exploration Mapping with Robot Semantic Mapping Semantic Updating
-SemanticFusion [13] × × ×
-DA-RNN [25] × × ×
-Hydra[31] × × ×
-ConceptGraphs[14] × × ×
-Dengler et al.[12] × ×
-Khronos[32] ×
-Ours
-TABLE I: Comparison of methods for autonomous exploration, semantic mapping and updating (reflecting environment changes).
-Figure 2: Illustration of our system for autonomous exploration, semantic mapping and map update.
 III System
 Our system consists of two phases.
-In the initial phase, the mobile robot explores the environment and builds the occupancy map incrementally.
-In the subsequent phase, the robot leverages the built occupancy map to accurately localize itself within the environment and proceeds to construct the semantic map.
+In the initial phase, a mobile robot explores the environment and builds the occupancy map incrementally.
+In the second phase, the robot leverages the built occupancy map to accurately localize within the environment and performs semantic mapping.
 The motivation to separate the process into two stages, rather than performing simultaneously, is because of computational efficiency and robustness.
-Attempting to perform both tasks in a single stage could lead to resource constraints and reduced performance, especially for large-scale environments with a large number of objects.
+Performing both tasks in a single stage leads to resource constraints and reduced performance, especially in large-scale environments with many objects.
 Additionally, in large environments, as the number of detected objects increases, the number of map corrections (like loop closures) applied to the detected objects increase, which slows down the mapping process.
-Evidently, this has also been mentioned in [32].
-Therefore, our two-phase approach is more efficient for semantic exploration and mapping of large-scale environments.
-An overview of the system is presented in Fig.
+This has also been mentioned in [26].
+Therefore, our two-phase approach is more suitable for semantic mapping of large-scale environments.
+Furthermore, this separation provides a significant advantage for Lidar-based mapping that relies on scan matching, such as GMapping [35] and Hector SLAM [36].
+Robot poses are not reliably maintained during the initial mapping for these methods as they rely on scan-matching.
+Accurate robot poses are obtained during the localization within a built map.
+Our two-stage approach allows these techniques to benefit from a reliable robot pose and thus more reliable object poses during the semantic mapping stage.
+An overview of our system is presented in Fig.
 2.
@@
 We have developed an environment exploration method based on the popular frontier exploration module [37], using a dynamic search window to efficiently navigate to frontiers in large real-world environments.
-When using [35] in large scale environments, as the map size increases, far-away frontiers become insignificant due to high cost, thus leaving a part of the map unexplored.
-Moreover, tuning the cost parameters in [35] is relatively hard and not intuitive.
-Figure 3: Environment traversal trajectory planning.
-Dynamic Window Frontier Exploration.
-We modified the frontier exploration module [35] by dynamically adjusting the robot’s search area between local and global regions [^2] Code available at https://github.com/IRVLUTD/dynamic-window-frontier-exploration.
+When using the original approach in large-scale environments, as the map size increases, far-away frontiers become insignificant due to high cost, thus leaving a part of the map unexplored.
+Moreover, tuning the cost parameters in the original method is not intuitive for large environments.
+Dynamic Window Frontier Exploration (DWFE).
+We modified the original frontier exploration method  [37] by dynamically adjusting the robot’s search area between local and global regions, as detailed in Algorithm 1222Code available at https://github.com/IRVLUTD/dynamic-window-frontier-exploration.
 This method identifies frontiers in the current search area and the robot navigates towards them to progressively map the environment.
 The search window is adjusted dynamically based on the density of the frontiers available around the robot.
-If the number of frontiers within the local search radius around the robot falls below a threshold, the search expands to a global radius.
+If the number of frontiers within the local search radius around the robot falls below a threshold, the search expands to a global radius, as seen in Fig.
+3.
 Setting this parameter to a large value expands the search region to the entire map.
-This allows the robot to move progressively ahead, reducing redundant coverage of the environment.
+This allows the robot to move progressively ahead, covering the local regions well, reducing redundant coverage of the environment.
 The exploration continues until the number of frontiers falls below a global threshold or after a fixed time has elapsed, at which point the exploration terminates and saves the built map.
@@
 2(a).
-During this process, the robot exploration trajectory is recorded which is later used to obtain a traversal plan.
+During this process, the robot exploration trajectory is recorded, which is later used to obtain a traversal plan.
 The saved occupancy map is then utilized to localize the robot during the semantic map construction and update phases.
-III-B Autonomous Construction of Semantic Map
-We utilize the saved occupancy map and the recorded robot exploration path to plan a trajectory to traverse the whole environment to construct the semantic map.
+Algorithm 1 Dynamic Window Frontier Exploration
+1:Input: Occupancy map $M$, local, global radii $r_{l}$, $r_{g}$, local, global threshold $T_{l}$, $T_{g}$, max time $T_{max}$, weights $\alpha,\beta$
+2:Initialize robot position $p$, time $t\leftarrow 0$
+3:while $t<T_{max}$ do
+4:  $F\leftarrow$ frontiers within $r_{l}$ around $p$
+5:  if $|F|<T_{l}$ then
+6:   $F\leftarrow$ frontiers within $r_{g}$ around $p$
+7:   if $|F|<T_{g}$ then
+8:     break $\triangleright$ Terminate if not enough frontiers
+9:   end if
+10:  end if
+11:  $f^{*}\leftarrow\arg\min_{f\in F}\left(\alpha\cdot distance(p,f)+\beta\cdot\text{size}(f)\right)$
+12:  yield $f^{*}$ $\triangleright$ Output the best frontier for this iteration
+13:  Navigate to $f^{*}$
+14:  $t\leftarrow t+1$
+15:end while
+Figure 3: Our dynamic search window during exploration.
+Figure 4: Environment traversal trajectory planning.
+III-B Autonomous Semantic Mapping
+We utilize the saved occupancy map and the recorded robot exploration path to plan a trajectory to traverse the environment again swiftly, to construct the semantic map.
 III-B1 Environment Traversing
@@
 To optimize this travel, we first sample a subset of $n$ points $P=\{p_{1},p_{2},\dots,p_{n}\}$ along the robot exploration path (Fig.
-3(a)), with a minimum of $2m$ separation between each pair of points (Fig.
-3(b)).
+4(a)), with a minimum of $2m$ separation between each pair of points (Fig.
+4(b)).
 We then construct a graph $G(P,E)$ with the sampled way points as nodes and the Euclidean distance $d$ between them as corresponding edge weights $w_{ij}=d_{ij}=\|p_{i}-p_{j}\|$.
 This is depicted in Fig.
-3(c).
+4(c).
 Finally, a traversal trajectory $T_{s}$ is obtained from $P$ using a greedy formulation of the Traveling Salesmen Problem (TSP) as explained in Algorithm 2.
 The generated trajectory is shown in Fig.
-3(d).
+4(d).
 Algorithm 2 Greedy Algorithm for TSP
@@
 4:Initialize visited $\leftarrow\{p_{1}\}$
-5:for $t\leftarrow 1$ to $n-1$ do
+5:for $t\leftarrow 1$ to n1n-1 do
 6:  Let $p_{\text{last}}\leftarrow$ last waypoint in the $T_{s}$
@@
 III-B2 Object Detection and Segmentation
-During the traversal, object detection and segmentation is performed in real time to construct the semantic map.
-For real-time object detection, the system leverages GroundingDINO [33] to detect objects in the environment.
-It takes in the current RGB image from the robot’s camera, identifies objects, and generates labels and bounding boxes for each detection.
-These bounding boxes are used as prompts for MobileSAM [38], a faster version of the Segment Anything Model (SAM) [39], to generate the object segmentation masks.
-After obtaining object segments of the current observation and its corresponding depth image, we compute the 3D point cloud $\mathbf{P}_{\text{camera}}$ for each segment in camera frame by applying camera intrinsic parameters to the depth.
-These points are then transformed to the map frame of reference through the transformation matrices representing camera pose in robot base frame, $\mathbf{T}_{\text{camera}}\in\mathbb{SE}(3)$ and robot pose in map frame $\mathbf{T}_{\text{robot}}\in\mathbb{SE}(3)$:
-$\mathbf{P}_{\text{map}}=\mathbf{T}_{\text{robot}}\cdot\mathbf{T}_{\text{camera}}\cdot\mathbf{P}_{\text{camera}}.$ (1)
-Finally, the mean position of the point cloud is computed using $\mathbf{P}_{\text{mean}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{P}_{\text{map}}(i)$ for $m$ points in the point cloud as the 3D position of the corresponding object in the map frame.
-III-B3 Semantic Map Representation and Construction
+During traversal, object detection and segmentation are performed in real time to build the semantic map.
+Our system utilizes GroundingDINO [33] to detect objects from the robot’s RGB image observation, providing labels and bounding boxes, which are then used as prompts for MobileSAM [39], a faster version of SAM [40], to generate segmentation masks.
+After obtaining object segments, we compute the 3D point cloud $\mathbf{P}_{\text{camera}}$ for each segment in the camera frame by applying the intrinsic parameters of the camera to the depth image observation.
+These points are then transformed into the map frame $M$, through the transformation matrices representing the camera pose in the robot base frame, $\mathbf{T}_{\text{camera}}\in\mathbb{SE}(3)$ and the robot pose in the map frame $\mathbf{T}_{\text{robot}}\in\mathbb{SE}(3)$, as follows: $\mathbf{P}_{\text{map}}=\mathbf{T}_{\text{robot}}\cdot\mathbf{T}_{\text{camera}}\cdot\mathbf{P}_{\text{camera}}$.
+Finally, the mean position of the point cloud is computed and represented as the 3D position of the corresponding object in the map frame.
+III-B3 Semantics Representation and Mapping
 Our semantic map is a hybrid representation.
 At the lower level, we utilize a 2D occupancy grid map $M$ that captures the spatial information of the free space and obstacles in the environment.
-At the higher level, we construct a topological map $G_{T}(V)$, where $V=\{v_{1},v_{2},\dots,v_{n}\}$ are the nodes representing uniquely detected objects (Fig.
+At the higher level, we construct a topological map defined by a set of vertices $V=\{v_{1},v_{2},\dots,v_{n}\}$.
+Each vertex $v_{i}$ represents a uniquely detected objects (Fig.
 2(d)).
@@
   • $v_{i}.\text{confidence}$: the confidence score of the detected object from object detection
-This representation facilitates fast object association and easy update of the nodes in the event of changes in the environment.
-Construction.
-To construct the semantic map, we first load the occupancy map $M$ built in Section III-A and localize the robot using it.
-Then an empty topological map $G_{T}(V)$, $V=\{\}$ is initialized.
-Additionally, for each object category (e.g., tables, chairs, doors), a category-specific object association distance threshold $\delta_{\{\text{category}\}}$ is defined.
-The robot then starts tracking the traversal trajectory generated in Section III-B1.
-Simultaneously, object detection and segmentation module is executed to detect objects in the field of view of the robot and compute their positions in the map $M$ frame of reference as illustrated in Fig.
-4(a).
-As depicted in Fig.
-4(b), for every detected object in the robot field of view, its minimum distance to the existing nodes of the corresponding category is calculated.
-If this distance is small than $\delta_{\{\text{category}\}}$, then that particular object is considered to be already represented on the map and hence not added to the topological map $G_{T}$ again.
-If the minimum distance is greater than the threshold, then the object is added as a new node to $G_{T}$ as illustrated in Fig.
-4(b).
-The result of this partial construction step is shown in Fig.
-4(c).
-This process is continued until the robot completes the tracking of the traversal trajectory, with new nodes being added to $G_{T}$ after object association.
-In this way, the semantic map is fully constructed, combining the geometric structure of the occupancy map with the layered semantics of the topological map.
-Figure 4: Adding objects to semantic map.
-Updating.
-Once some changes are made to the environment, we update the semantic map to reflect these changes.
-This process is similar to the construction phase, but with additional validation steps to handle existing object removal, displacement, and new object addition.
-First, the constructed semantic map is loaded, which acts as a reference for the current state of the environment.
-As the robot tracks the traversal trajectory again, the detected objects in its current field of view are compared against those objects that are expected to be in the same field of view in the semantic map.
-If an expected object is no longer detected, it is assumed to have been removed or moved and the corresponding node is deleted from the map as shown in Fig.
-5.
-Similarly, newly detected objects that do not match any existing nodes within the distance threshold $\delta_{\{\text{category}\}}$ are added as new nodes, representing new or relocated objects.
-This process updates the semantic map in real-time.
-Figure 5: Removing objects from semantic map.
+This representation facilitates fast object association and easy update of nodes during environment changes.
+Semantic Mapping and Updating.
+To build the semantic map, we first load the occupancy map $M$ built as detailed in Sec.
+III-A and localize the robot within it.
+If the semantic mapping process is performed for the first time, a topological map is initialized with an empty set of vertex $V^{\prime}=\emptyset$.
+Otherwise, the previously constructed semantic map is loaded ( $V^{\prime}\leftarrow V$).
+For each object category (e.g., tables, chairs, doors), a category-specific object association distance threshold $\delta_{\{\text{category}\}}$ is defined to determine whether a detected object corresponds to an existing one.
+Categories without explicit thresholds are assigned a default value.
+In practice, setting $\delta$ roughly equal to the typical object size improves the differentiation between nearby objects.
+Algorithm 3 Semantic Mapping and Updating
+1:Input: Traversal trajectory $T_{s}$, Distance thresholds $\Delta=\{\delta_{c}\mid c\in\text{categories}\}$
+2:Output: Updated semantic map $G_{T}(V^{\prime})$
+3:if no prior map exists then
+4:  Initialize $V^{\prime}\leftarrow\emptyset$, id $\leftarrow 0$
+5:else
+6:  Load $G_{T}(V)$
+7:  $V^{\prime}\leftarrow V$,  $\text{id}\leftarrow\max_{v\in V^{\prime}}v.\text{id}+1$
+8:end if
+9:while Traversing $T_{s}$ do
+10:  $\mathcal{O}_{\text{current}}\leftarrow$ objects detected in FoV
+11:  $\mathcal{V}_{\text{expected}}\leftarrow$ expected nodes in FoV from $G_{T}(V^{\prime})$
+12:  $\mathcal{O}_{\text{unmatched}}\leftarrow\mathcal{O}_{\text{current}}$ $\triangleright$ Initialize unmatched objects
+13:  for $v\in\mathcal{V}_{\text{expected}}$ do
+14:   if no $o\in\mathcal{O}_{\text{current}}$ such that
+[…]
+[…]
+17:     $V^{\prime}\leftarrow V^{\prime}\setminus\{v\}$ $\triangleright$ Remove node
+18:   else
+19:     $\mathcal{O}_{\text{unmatched}}\leftarrow\mathcal{O}_{\text{unmatched}}\setminus\{o\}$
+20:   end if
+21:  end for
+22:  for $o\in\mathcal{O}_{\text{unmatched}}$ do
+23:   $V_{c}\leftarrow\{v\in V^{\prime}\mid v.\text{category}=o.\text{category}\}$
+24:   $d_{\min}\leftarrow\infty$
+25:   if $V_{c}\neq\emptyset$ then
+26:     $d_{\min}\leftarrow\min_{v\in V_{c}}\text{dist}(P_{\text{mean}}(o),v.\text{position})$
+27:   end if
+28:   if $d_{\min}>\delta_{o.\text{category}}$ then
+29:     $v_{new}=(\text{id},o.\text{category},$
+30:          $P_{\text{mean}}(o.\text{segment}),$
+31:          $o.\text{confidence})$
+32:     $V^{\prime}\leftarrow V^{\prime}\cup\{v_{new}\}$ $\triangleright$ Add new node to $V^{\prime}$
+33:     $\text{id}\leftarrow\text{id}+1$
+34:   end if
+35:  end for
+36:end while
+37:Return $G_{T}(V^{\prime})$
+Once initialized, the robot then begins following the traversal trajectory $T_{s}$ (see Sect.
+III-B1).
+Simultaneously, the object detection and segmentation module processes the current image observation of the robot’s field of view (FoV) to detect objects $\mathcal{O}_{\text{current}}$ and compute their positions in the map $M$ frame of reference as seen in Fig.
+5(a).
+These detections are initially considered unmatched $\mathcal{O}_{\text{unmatched}}$.
+At each step, an approximate polygon representing the field of view is derived from the limits of the current depth point cloud.
+Using this FoV polygon, the semantic map is queried to retrieve the nodes $\mathcal{V}_{\text{expected}}$, which represent the objects expected to be within the FoV.
+The Semantic Mapping procedure (described in Alg.
+3 iteratively updates the vertex set $V^{\prime}$.
+Node Removal.
+For each expected node $v\in\mathcal{V}_{\text{expected}}$, the algorithm checks for a matching object $o\in\mathcal{O}_{\text{current}}$ of the same category ($o.\text{category}=v.\text{category}$) and within the threshold distance $\text{dist}(P_{\text{mean}}(o),v.\text{position})\leq\delta_{v.\text{category}}$.
+If no match is found, the node $v$ is removed from $V^{\prime}$, indicating that the corresponding object is no longer present in the expected location (e.g., moved or removed).
+The matched objects are removed from $\mathcal{O}_{\text{unmatched}}$.
+Node Addition.
+For each unmatched object $o\in\mathcal{O}_{\text{unmatched}}$, the algorithm identifies nodes $V_{c}=\{v\in V^{\prime}\mid v.\text{category}=o.\text{category}\}$ belonging to the same category.
+If no such category exists $V_{c}=\emptyset$ or an object is not associated with existing ones $\min_{v\in V_{c}}\text{dist}(P_{\text{mean}}(o),v.\text{position})>\delta_{o.\text{category}}$, a new node $v_{new}=(\text{id},o.\text{category},P_{\text{mean}}(o.\text{segment}),o.\text{confidence})$ is added to $V^{\prime}$.
+The object association mechanism is illustrated in Fig.
+5(b).
+This iterative process continues until the robot completes the traversal of $T_{s}$.
+By continuously updating the vertex set $V^{\prime}$, the algorithm ensures that the semantic map captures both short-term and long-term environmental changes.
+The resulting map seamlessly integrates the geometric structure of the occupancy map with the semantic information of the topological map , enabling robust and faster scene understanding.
+Figure 5: Updating Semantic Map through Object Association.
 IV Experiments
-We have conducted experiments in two different environments with different object categories and evaluate the performance of our system in those environments.
+To evaluate the effectiveness of our proposed Dynamic Window Frontier Exploration method (Alg.
+1), we conducted experiments in a large-scale simulated hospital environment[41] of approximately $1440~\text{sq.m}$, in the Gazebo simulator.
+We conducted 20 iterations of exploration, where in each iteration a Fetch mobile manipulator robot was spawned at a different starting location and explored for 15 minutes.
+The same starting locations are used for both our approach and conventional frontier exploration.
+During each run, we recorded the area covered and the path length traveled, allowing a direct comparison between our approach and the conventional frontier exploration baseline [37].
+[…]
+Baseline[37] 669.12 107.10
+DFWE (ours) 721.05 94.30
+TABLE II: Average area coverage and path length for baseline vs. DFWE in a large-scale simulated environment.
+The results in Table II demonstrate the improved exploration efficiency using our method.
+Compared to the baseline exploration, our method explores $\mathbf{\approx 7.8\%}$ more area while traveling $\mathbf{\approx 12\%}$ less on average.
+These numbers are expected to get better as the size of environment increases.
+We conducted semantic mapping experiments in two different real-world environments with different object categories and evaluated the performance of our system.
 Environment A.
-The fourth floor of the ECSS building at UT Dallas is a large indoor environment measuring approximately $93m\times 90m$.
-It consists of 18 corridors with a total traversal length of approximately $800m$.
-These corridors were mostly furnished with typical office furniture, including doors, tables, and chairs.
-Therefore, we only consider these 3 classes, i.e., table, door and chair for semantic mapping of this environment.
+This is a large indoor environment measuring approximately $\mathbf{93m\times 90m}$.
+It consists of 18 corridors with a total traversal length of approximately 800 m These corridors were mostly furnished with typical office furniture, including doors, tables, and chairs.
+Therefore, we only consider these 3 classes, i.e., table, door and chair for reliable semantic mapping of this environment.
 Environment B.
-This environment is a laboratory environment with a reduced size measuring $9m\times 13m$.
+This is a laboratory environment with a reduced size measuring $9m\times 13m$.
 This space included a more diverse set of objects such as televisions, trash bins, umbrellas, chairs, cabinets and persons.
-Robot. We run the experiments on a Fetch mobile robot equipped with a LiDAR sensor and an RGB-D camera as depicted in Fig. 6.
-Figure 6: The Fetch robot maps the environment using a $220^{o}$ laser scanner of $25m$ range, while leveraging depth observations for obstacle avoidance.
-Its RGB-D camera operates at 30fps, enabling real-time object detection and segmentation for semantic mapping running on the connected GPU laptop.
-Exploration and Mapping.
-To autonomously explore and map the environments, we integrate the mapping, exploration, and navigation modules.
-The Gmapping ROS package [37] was used to incrementally build a 0.1m/pixel occupancy grid map, using laser scan data from the Fetch robot.
-Simultaneously, we ran our custom-built Dynamic Window Frontier Exploration module to identify and select best valued frontier in the search window.
-The move_base ROS package [40] guided the robot to these frontiers by combining a global planner for long term path planning and a local planner that dynamically avoids obstacles not seen during global planning.
-This ensures safe navigation when any new objects are added or moved around in the environment.
-The robot’s maximum velocity was limited to 0.6m/s, with a 0.7m inflation radius to prevent it from traveling too close to the obstacles.
-In Environment A, after 150 minutes of exploration, which also included several stops due to close proximity of persons in narrow corridors, the environment was fully mapped, and 2,250 trajectory points were recorded at 0.25Hz.
-In Environment B, exploration was completed in approximately 4 minutes, with 212 points recorded at 1Hz.
+Only objects with a detection confidence of 0.6 or higher are considered.
+Robot. We run the experiments on a Fetch mobile robot equipped with a LiDAR sensor and an RGB-D camera.
+IV-A Real World Exploration and Mapping
+We integrate mapping, exploration, and navigation to autonomously map environments.
+Using Gmapping [38], a 0.1 m/pixel resolution occupancy map was built using laser scans, while our Dynamic Window Frontier Exploration module selected the best frontiers in the search radius.
+The move_base ROS package [42] guided the robot safely to these frontiers with global and local planning, dynamically avoiding unexpected obstacles.
+The robot was limited to 0.6 m/s speed with a 0.7 m inflation radius for safe navigation in the environment.
+In Environment A, 150 min of the exploration (including stops for people in narrow corridors) produced a full map and 2,250 trajectory points in 0.25 Hz.
+In Environment B, the exploration was completed in approximately 4 min, recording 212 points at 1 Hz.
 Environment Traversing.
-The greedy TSP described in Sect.
-III-B1 is formulated using the poses recorded during the exploration, resulting in traversal trajectories of 130 poses for Environment A and 15 poses for Environment B.
-In particular, the time to revisit is significantly reduced compared to the exploration phase.
-Revisiting took approximately 35 minutes for Environment A and 2 minutes for Environment B.
-This highlights the efficiency gained through this traversal optimization.
-Real-time Construction and Updating of Semantic Maps.
-To enable real-time semantic mapping, the Fetch robot was equipped with a laptop featuring NVIDIA RTX 4090 GPU, running Ubuntu 20.04 and ROS Noetic.
-The laptop and Fetch robot were connected via Ethernet for subscribing and publishing the data through ROS topics.
-Initially, the environment remained unchanged compared to its state during exploration.
+Using the greedy TSP from Sec.
+III-B1, traversal trajectories were generated from the recorded exploration poses: 130 poses for Environment A and 15 poses for Environment B.
+Revisiting was significantly faster, taking approximately 35 min for Environment A and 2 min for Environment B, demonstrating the efficiency of the optimized traversal.
+IV-B Real-time Semantic Mapping
+For real-time semantic mapping, we equip Fetch with a laptop with an RTX 4090 GPU running Ubuntu 20.04 and ROS Noetic.
 The system then begins with the built occupancy map.
-After localizing itself using AMCL ROS [40], the robot progressively builds the semantic map.
-Subsequently, semantic map updates were tested by removal, addition, and/or moving the objects in the environments.
-Visual results of environment A can be found in Fig.
-2.
-More details of it can be found in the supplementary video.
+After localizing using AMCL ROS [42], the robot progressively builds the semantic map.
+We set the object association distance threshold of 0.7 m for all categories.
+Subsequently, two additional iterations are performed, during which the environment is modified through the removal, addition, and/or relocating objects in the environments.
 Fig.
-7 depicts some of the changes in Environment B, reflected in the semantic updating stages.
-Figure 7: Visualization of some selected environmental changes in Environment B across two update stages.
-In Update stage I, 2 trash bins are removed.
-After that in update stage II, 2 chairs were removed.
+6 shows some of the changes in Environment B, reflected during semantic updating.
+Env Class Iteration I (Initial run) Iteration II (vs I) Iteration III (vs I)
+   GT Map $\DeltaGT Δ\DeltaMap$ $\DeltaGT Δ\DeltaMap$
+A Table 45 45 0 -6 0* -8
+Chair 109 84 -10 -11 0* -15
+Door 170 153 0 +2 0 +3
+B Person 2 2 0* 0 0 0
+Trash Bin 2 2 -2 -2 -2 -1
+Umbrella 1 1 -1 -1 -1 -1
+Chair 3 3 0 0 -2 -2
+Television 1 2 0 -1 0 -1
+Cabinet 3 2 0 -1 0 0
+[…]
+Figure 6: Visualization of environmental changes in Environment B: Iteration II (2 trash bins removed) and Iteration III (2 chairs removed).
 When robot revisited, these changes are reflected in the updated semantic map.
-Env Class Construction Update I Update II
-GT Map GT Map GT Map
-A Table 45 45 45 39 45* 37
-Chair 109 84 99 73 109* 69
-Door 170 153 170 155 170 156
-B Person 2 2 2* 2 2 2
-Trash Bin 2 2 0 0 0 1
-Umbrella 1 1 0 0 0 0
-Chair 3 3 3 3 1* 1
-Television 1 2 1 1 1 1
-Cabinet 3 2 3 1 3 2
-TABLE II: Semantic map object count across the construction and update stages for A & B Environments.
-GT: Ground Truth, Map: no.
-of objects mapped.
-The table compares the ground truth with the actual results across the construction, update I, and update II scenarios.
-* indicates that some objects have been relocated from their earlier position.
-IV-A Quantitative Analysis
-Table II summarizes the number of objects detected in both Environment A and Environment B during the Construction, Update I and Update II phases.
-In Environment A, we can see that while doors and tables were stably detected, the detection of chairs was relatively more challenging.
-During update I, the removal of 10 chairs after the construction phase is reflected in the decrease in the number of chairs detected from 84 to 73.
-However, in Update II, only 69 chairs were detected, with false negatives increasing to 29 as shown in Table III.
-This was mainly because some relocated chairs were placed in the path of the robot traversal trajectory.
-As a result, robot took a detour to avoid hitting them, causing it to miss some chairs that were not relocated and previously detected.
-Moreover, people occupying chairs occluded them, increasing false negatives.
-In contrast, Environment B has very stable detections, with many classes having 100% precision.
-However, some categories like cabinets that are dark colored, are challenging to detect in nominal lightning, resulting in them not being detected.
-IV-B Qualitative Analysis
+IV-C Quantitative Analysis.
+Table III reports the semantic mapping results across three iterations, with Iteration I serving as the reference.
+For Iteration I, we show the absolute ground truth (GT) objects and those mapped by the robot (Map).
+For Iterations II and III, we report only the relative changes ($\DeltaGT, Δ\DeltaMap$) compared to Iteration I.
+[…]
+[…]
+Overall, this analysis shows that while stable objects (e.g., tables, doors) maintain consistent mapping across iterations, dynamic or occlusion-prone categories such as chairs exhibit greater variability in detection and mapping performance.
+Class Iteration I Iteration II Iteration III
+TP FP FN F1 TP FP FN F1 TP FP FN F1
+Environment A
+Table 36 9 9 0.8 36 3 9 0.86 33 4 12 0.80
+Chair 84 0 25 0.87 72 1 27 0.84 67 2 42 0.75
+Door 140 13 30 0.87 140 15 30 0.86 142 14 28 0.87
+Environment B
+Person 2 0 0 1.00 2 0 0 1.00 2 0 0 1.00
+Trash Bin 2 0 0 1.00 0 0 0 - 0 1 0 0.00
+Umbrella 1 0 0 1.00 0 0 0 - 0 0 0 -
+Chair 3 0 0 1.00 3 0 0 1.00 1 0 0 1.00
+Television 1 1 0 0.67 1 0 0 1.00 1 0 0 1.00
+Cabinet 2 0 1 0.80 1 0 2 0.50 2 0 1 0.80
+TABLE IV: Semantic map object count across Iteration I, II & III for A & B Environments. TP: True Positives, FP: False Positives, FN: False Negatives, F1 scores are reported.
+Figure 7: Intermediate semantic maps generated by : (a) our method, (b) Khronos [26], running at 21FPS of data replay.
+We compare and evaluate our semantic mapping approach with Khronos [26] on the real-world mezzanine scene dataset [43] provided as a ROS bag.
+Fig.
+7 represents the semantic maps generated by both methods at an intermediate stage.
+The results are summarized in Table V.
+At 30 FPS of ROS bag replay, our method successfully maps 79 objects, while Khronos only maps 3, showing that our approach can process visual information faster and efficient as well.
+When the input rate is reduced to 21 FPS, Khronos maps 87 objects, while our method achieves a comparable 86.
+Data Replay Rate Objects Mapped by Method
+Khronos [26] Ours
+21 FPS 87 86
+30 FPS 3 79
+TABLE V: Number of objects mapped by each method at different data FPS on the real-world mezzanine dataset.
+This performance gap arises from the fundamental differences in system design.
+Since Khronos performs SLAM and semantic mapping in a single stage, its runtime is to an extent governed by loop-closure optimization, achieving superior SLAM performance, but limits the scalability of semantic mapping at higher frame rates.
+In contrast, our method decouples occupancy and semantic mapping, allowing it to achieve both higher throughput and competitive accuracy across different operating conditions.
+IV-D Qualitative Analysis
 Figure 8: Examples of accurate detection and segmentation of chairs, tables and doors in Environment A.
-Class Construction Update I Update II
-TP FP FN Precision TP FP FN Precision TP FP FN Precision
-Environment A
-Table 36 9 4 0.9 36 3 9 0.8 33 4 6 0.85
-Chair 84 0 16 0.84 72 1 23 0.76 67 2 29 0.69
-Door 140 13 28 0.83 140 15 26 0.84 142 14 28 0.84
-Environment B
-Person 2 0 0 1 2 0 0 1 2 0 0 1
-Trash Bin 2 0 0 1 0 0 0 - 0 1 0 0
-Umbrella 1 0 0 1 0 0 0 - 0 0 0 -
-Chair 3 0 0 1 3 0 0 1 1 0 0 1
-Television 1 1 0 0.5 1 0 0 1 1 0 0 1
-Cabinet 2 0 1 0.67 1 0 2 0.33 2 0 1 0.67
-TABLE III: Semantic map object count across the construction and update stages for A & B Environments.
-TP: True Positives, FP: False Positives, FN: False Negatives, Precision: Precision for detected objects across the construction and update stages.
+Fig. 8 illustrates several examples of accurate object detection during our experiments for chairs, tables, and doors.
 In Fig. 9, we show several challenging scenarios in object detection and segmentation, while robot is navigating.
@@
 Several of such doors are not detected, leading to a reduction in the number of object nodes in the semantic map.
-This is evident in Table II.
+This is evident in Table III.
 Figure 9: Representation of various challenging scenarios involved in object detection and semantic mapping.
@@
 This representation enables us to update the semantic map efficiently to reflect the changes in the environment.
-Experiments conducted on a mobile manipulatory system in a large- and medium-scale indoor environments demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
+Experiments conducted on a mobile manipulatory system in large- and medium-scale indoor environments demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
 Object perception during robot navigation still remains a challenge.
@@
 Acknowledgment
-This work was supported by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005 and the Sony Research Award Program.
+This work was supported by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005, the Sony Research Award Program, the National Science Foundation (NSF) under Grant Nos.
+2346528 and 2520553, and the NVIDIA Academic Grant Program Award.
 The work of T.